OpenStudy (anonymous):

solve: lim(x->0) of (1 - cosx + 3sinx)/x

OpenStudy (anonymous):

\[ \lim_{x\to 0}\frac{1-\cos x+3\sin x}{x} = -\lim_{x\to 0}\frac{\cos x-1}{x}+3\lim_{x\to 0}\frac{\sin x}{x} \]

OpenStudy (anonymous):

These two limits are worth remembering.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is the next step

OpenStudy (anonymous):

Look up those limits

OpenStudy (anonymous):

the answer is 3

OpenStudy (anonymous):

other than knowing those two limits, are there any other ways to solve (aside from finding those limits by other means)

OpenStudy (freckles):

wio has the most simplest way but you could use l'hospital since 1-cos(0)+3sin(0)=1-1+0=0 and 0=0 we have the case 0/0 So we can apply the l'hospital: \[\lim_{x \rightarrow 0}\frac{1-\cos(x)+3\sin(x)}{x}=\lim_{x \rightarrow 0}\frac{0+\sin(x)+3\cos(x)}{1}= \lim_{x \rightarrow 0}(\sin(x)+3\cos(x)) \\ =\sin(0)+3\cos(0) =0+3(1)=0+3=3\]